# Collaborative Perception Datasets for Autonomous Driving: A Review

### v1

@@ I Introduction
Perception serves as a foundational component of Autonomous Driving (AD) systems [1], [2], [3], [4], [5], [6], [7], [8], [9], enabling the vehicle to perceive its surroundings through various onboard sensors and to perform real-time tasks such as object detection [10][11], semantic segmentation, tracking, and trajectory prediction in dynamic traffic environments [12]. The precision and reliability of the perception system directly impact the safety and decision-making capability of autonomous vehicles, making it a critical aspect to achieve high-level AD. In recent years, significant progress has been made in single-vehicle perception, driven by progress in computer vision [13][14][15], deep learning [16][17], and sensor technology [18][19]. Numerous studies have leveraged cameras [20][21][22], LiDAR [23][24], Radar [25][26], and multi-sensor fusion [27][28][29][30][31][32] to enhance environmental perception in AD systems. However, traditional single-vehicle perception is constrained by limited fields of view, sensor blind spots, and environmental occlusions, making it insufficient to meet the high-reliability requirements of AD in complex and dynamic traffic environments.
@@
To overcome the limitations of single-vehicle perception, Collaborative Perception (CP) leverages Vehicle-to-Everything (V2X) communication to enable real-time information sharing among vehicles and infrastructure [33], thereby enhancing the perception of occluded and long-range regions, improving driving safety, and expanding the Operational Design Domain (ODD) [34][35]. For instance, at an urban intersection, a single vehicle may fail to detect pedestrians or oncoming vehicles that are occluded by buildings or other obstacles. By leveraging V2X connectivity, autonomous vehicles can receive supplementary perception data from roadside infrastructure or neighboring vehicles, thereby improving situational awareness and reducing accident risk. V2X communication encompasses Vehicle-to-Vehicle (V2V), Vehicle-to-Infrastructure (V2I), Infrastructure-to-Infrastructure (I2I), and Vehicle-to-Network (V2N) links, empowering AD systems to overcome the inherent limitations of individual perception and achieve a more holistic understanding of their surroundings.
High-quality datasets are essential for advancing algorithm development and performance evaluation in CP research [36]. Existing single-vehicle perception datasets, such as KITTI [37], nuScenes [38], and Waymo [2], are primarily designed for individual perception tasks and do not support multi-vehicle or infrastructure-level information exchange, rendering them unsuitable for direct use in CP studies.
In recent years, numerous CP datasets, as illustrated in Fig. 1, have been released to support various scenarios, tasks, and sensor modalities, including DAIR-V2X [39], OPV2V [40], V2X-Sim [41], and V2X-Seq [42]. While these datasets offer valuable resources for the CP community, they differ significantly in terms of scale, scene diversity, annotation quality, sensor configurations, and the simulation of communication latency. A systematic review and analysis of existing CP datasets can help researchers better understand their characteristics, application boundaries, and limitations, thereby providing scientific guidance for future investigations. Furthermore, identifying existing gaps and prospective research directions will contribute to the continued advancement of CP technologies.
@@ I-A Roadmap
Figure 2: Chronological overview of collaborative perception datasets, including roadside perception, V2V, V2I, V2X, and I2I datasets.
Fig. 2 illustrates the chronological evolution of CP datasets. Overall, the development of these datasets has progressed through several key phases, each marking a significant shift in perception paradigms or improvements in dataset diversity. Changes in collaboration models, increasing scene complexity, the expansion of supported tasks, and the continuous growth in sensor configurations have all contributed to the advancement of CP technologies.
In the early stages, datasets were primarily focused on infrastructure-based perception, as exemplified by Ko-PER [43], CityFlow [44], and the INTERACTION [45]. However, these datasets did not support V2V or V2I information exchange, thereby limiting their applicability to CP research. With the advancement of V2X communication technologies, the number of related datasets has steadily increased. Around 2020, several V2V and V2I datasets were released, including T & J [46], V2V-Sim [47], CoopInf [48], and CODD [49], which collectively accelerated progress in CP research. However, datasets from this period still faced limitations in terms of scene complexity, sensor modality diversity, and task coverage. As the field progressed, newer CP datasets began to target a broader range of scenarios. For example, Adver-City [50] focuses on adverse weather conditions, DeepAccident [51] addresses accident prediction, and H-V2X [52] is tailored for highway environments. These datasets have significantly enhanced the adaptability of CP systems.
Since 2022, CP dataset development has increasingly emphasized multi-modal fusion and real-world deployment. Datasets such as V2V4Real [53], V2X-Real [54], and DAIR-V2X-C [39] incorporate real-world data collection to overcome the limitations of synthetic datasets in terms of authenticity and distribution. Simultaneously, newer datasets demonstrate broader task coverage and richer sensor modalities. For instance, V2X-Seq [42], V2XReID [55], and DAIR-V2X-Traj [56] support a variety of perception tasks, while V2X-Radar [57] and V2X-R [58] incorporate 4D radar to enhance perception quality. Moreover, the emergence of I2I datasets such as InScope [59] and Rcooper [60] has further enriched the collaborative paradigms within CP. In general, the evolution of CP datasets reflects a continuous trend toward greater complexity, richer modalities, and more specialized scenarios.
@@ I-B Comparison to Related Reviews
We aim to clarify the distinctions between this review and previous related reviews [61], [36], [62], [63], [64], [65], [66], [67], [68], [69], [70]. While some existing reviews partially overlap with ours in terms of dataset coverage [71] [72], they omit several important CP datasets and recent benchmark methodologies introduced in the past few years. Furthermore, their classification frameworks tend to be narrow and often lack systematic summarization and in-depth comparative analysis. In addition, some reviews concentrate primarily on vehicle-centric datasets [61][36][62][63], or explore specific aspects of CP such as multimodal fusion [64][65] or certain CP techniques [66][67][68][69][70], without offering a comprehensive overview of dataset evolution and their applicability across various scenarios. In contrast, this review presents the most up-to-date and complete collection of CP datasets, encompassing a broader spectrum of perspectives. It also delivers an in-depth discussion on dataset diversity, key challenges, and emerging research directions.
@@ I-C Main Contributions
In this review, we present a systematic and comprehensive analysis of CP datasets for AD, as summarized in Fig. 3. The main contributions of this work are as follows:
Comprehensive Review of CP Datasets: To the best of our knowledge, this is the first dedicated and comprehensive review focused specifically on CP datasets. We systematically categorize and analyze available datasets based on collaboration types, data sources, sensor modalities, and application scenarios.
Task-Oriented Dataset Classification: We classify CP datasets according to core perception tasks in AD, including 3D object detection, semantic segmentation, and object tracking and prediction. This task-centric taxonomy facilitates effective dataset selection based on research needs. In addition, we cover the latest developments in CP datasets, providing a timely and up-to-date overview.
Dataset Comparison and Emerging Trends: We conduct multi-dimensional comparisons, highlight high-impact datasets, and discuss key challenges and future directions such as standardization, LLM-driven annotation, and data diversity.
@@ I-D Review Structure
The structure of this paper is illustrated in Fig. 3. To better present our review, the remainder of this paper is organized as follows. Section II introduces datasets under different collaborative paradigms. Section III compares the characteristics of simulated and real-world datasets. Section IV reviews existing CP datasets from the perspective of different application scenarios. Section V analyzes datasets based on sensor modalities. Section VI summarizes and analyzes the CP tasks supported by these datasets, along with their corresponding evaluation metrics. Section VII provides a comprehensive analysis of existing CP datasets, focusing on their availability, quality, and impact. Section VIII discusses the challenges and trends in CP datasets. Section IX concludes the paper and outlines future research directions. A summary of all abbreviations used throughout the paper is provided in Table I.
@@
CAM Cooperative Awareness Message
CPM Cooperative Perception Message
CPS Collective Perception Service
CAD Collaborative Autonomous Driving
CDA Cooperative Driving Automation
GPS Global Positioning System
IMU Inertial Measurement Unit
@@ II Collaboration-Based Perception Datasets
With the rapid development of Intelligent Connected Vehicles (ICVs) and Intelligent Transportation Systems (ITS), CP paradigms have become increasingly diverse. These paradigms now encompass not only pure roadside perception but also V2V, V2I, and more comprehensive V2X. Additionally, emerging paradigms such as I2I collaboration have recently gained attention. As illustrated in Fig. 4, various CP architectures are now being explored. In this section, we classify and discuss CP datasets according to their communication paradigms, including roadside perception datasets (Section II-A), V2V datasets (Section II-B), V2I datasets (Section II-C), V2X datasets (Section II-D), and I2I datasets (Section II-E). Table II provides a comprehensive summary of representative datasets under each paradigm. These datasets play a crucial role in advancing cooperative intelligence for ICVs and ITS.
@@
SCOPE [93] 2024 arxiv C, L Sim 2DOD, 3DOD, SS, S2R 17k 17k 575k Link
@@ II-A Roadside Datasets
Roadside perception plays a vital role in AD and ITS by offering a global, occlusion-resilient view of traffic participants [94][95][96]. Unlike vehicle-mounted systems, infrastructure-based sensing units—typically installed on poles, gantries, or overpasses—enable the monitoring of complex intersections, dense traffic flows, and long-range environments with enhanced spatial stability and temporal continuity. By supplementing on-vehicle sensing with infrastructure-level observations, roadside perception significantly improves perception robustness in scenarios involving occlusions, blind spots, or congested conditions.
A variety of representative roadside datasets have been introduced to support infrastructure-based perception research. Early datasets such as Ko-PER [43] and CityFlow [44] primarily target intersection environments and person or vehicle Re-ID tasks. INTERACTION [45] expands the behavioral diversity of roadside scenarios by incorporating multi-national data with adversarial and cooperative driving maneuvers. Rope3D [75] provides high-resolution 3D data along with dedicated metrics for evaluating roadside 3D object detection. LUMPI [76] supports multi-view analysis using synchronized multi-sensor setups, while IPS300+ [74] offers dense, long-range 3D annotations for complex urban intersections. TUMTraf-I [77] enhances diversity through coverage of varying lighting and weather conditions. For highway-oriented research, A9-Dataset [73] offers real-world, high-speed traffic scenarios, whereas RoScenes [78] features the largest perception area to date through high-density, multi-camera roadside deployment. H-V2X [52] extends this trend by integrating radar and camera data, supporting a range of tasks including detection, tracking, and trajectory prediction under complex highway conditions.
The release of roadside datasets has significantly broadened the scope of perception research, enabling detailed studies on object detection, semantic segmentation, and multi-object tracking from an infrastructure-centric perspective. These datasets also provide a valuable foundation for future V2I CP systems by enhancing scene coverage and perceptual redundancy. However, several challenges remain inherent to infrastructure-based data collection. Such as, high deployment and maintenance costs restrict large-scale installations, particularly in highway and rural areas. Data distributions are often skewed toward urban environments, which limits the generalizability of trained models across diverse traffic scenarios. Additionally, precise temporal synchronization and spatial alignment among heterogeneous sensors remain non-trivial, especially in long-term or multi-site deployments.
@@ II-B V2V Datasets
V2V CP enhances perception range and accuracy by enabling information sharing among multiple vehicles. Compared to single-vehicle perception, V2V collaboration effectively mitigates blind spots caused by occlusions and compensates for the limitations of individual sensors in specific scenarios. This leads to improved robustness in tasks such as object detection, semantic understanding, and behavior prediction. Furthermore, the integration of multi-sensor data and cross-view information enables V2V systems to perform more adaptively in complex and dynamic environments. Therefore, the development of high-quality V2V CP datasets is of great importance for advancing research in multi-vehicle collaborative autonomous driving systems.
In recent years, V2V datasets have continued to evolve in terms of collaborative modeling and data organization mechanisms. Most datasets, including OPV2V [40], MARS [82], and LUCOOP [81], adopt an offline alignment strategy that unifies data from multiple vehicles into a common coordinate system using global positioning or timestamps. This approach simulates idealized synchronized collaboration scenarios. However, it fails to account for real-world issues such as asynchronous sampling, communication latency, and localization errors during actual V2V communication. To address these limitations, HEAL [83] introduces asynchronous perception and temporal alignment mechanisms, while OPV2V-N [97] employs neural modeling to improve perception robustness. COMAP [79] adopts a task-oriented modeling approach from a multi-agent SLAM perspective. In terms of agent configuration, most datasets support CP among 2 to 4 vehicles. MARS [82] enables data collection across multiple time periods, trajectories, and viewpoints using four connected vehicles. OPV2V+ [80] extends viewpoint variation modeling by supporting primary view selection and perception completion. Regarding task coverage, existing datasets support key tasks such as object detection, tracking, Bird’s Eye View (BEV) semantic segmentation, and path planning. V2V-QA [84] further incorporates multi-modal large language models to support intelligent question answering and path reasoning tasks. IRV2V [98] and CODD [49] focus on asynchronous robustness and multi-view point cloud registration, reflecting the expanding scope of V2V data modeling in terms of collaboration granularity and semantic capabilities.
Despite the notable progress made by V2V CP datasets in enabling multi-vehicle fusion research, several distinct challenges remain. For example, most existing datasets rely on offline stitching techniques for data integration and lack modeling of real communication links, transmission latency, and collaborative constraints. As a result, they fail to capture the uncertainty and dynamic characteristics of information exchange in real V2V scenarios, limiting deeper investigations into the essence of cooperative mechanisms. Looking ahead, the development of V2V datasets should focus on realistic modeling of communication processes, the design of dynamic alignment mechanisms for CP, and the construction of efficient multi-vehicle data acquisition and annotation pipelines in complex urban environments.
@@ II-C V2I Datasets
V2I perception datasets aim to enhance AD systems by enabling collaboration between vehicles and roadside infrastructure. These datasets incorporate multimodal data from both vehicle-mounted and infrastructure-based sensors, offering a more comprehensive and accurate understanding of the environment.
Representative V2I datasets exhibit distinct characteristics and collaborative strategies. CoopInf [48] employs eight infrastructure sensors for multi-view 3D object detection. DAIR-V2X-Seq [42] is the first large-scale sequential V2X dataset using data from both infrastructure and vehicles. HoloVIC [85] constructs holographic intersections with diverse layouts to overcome occlusion and blind spots through multisensor fusion. TUMTraf-V2X [87] emphasizes perception under complex lighting conditions and supports 3D detection and tracking. OTVIC [86] targets high-speed and noisy communication scenarios with multimodal, multi-view perception data.
While these datasets significantly contribute to improving V2I CP, challenges remain. Sensor synchronization is technically demanding, and many datasets offer limited scene diversity or weather coverage. Future efforts should prioritize building more diverse, comprehensive V2I datasets to improve perception robustness in complex and dynamic driving environments.

### v2

@@ I Introduction
Perception is fundamental to Autonomous Driving (AD) systems [1][2][3], [4][5][6][7][8], enabling vehicles to understand their surroundings via onboard sensors for tasks such as object detection [9][10], semantic segmentation, tracking [11], and trajectory prediction [12]. The accuracy and robustness of perception directly affect safety and decision-making. Recent advances in computer vision [13], [14], [15], deep learning [16][17], and sensor technology [18], [19][20] have led to substantial progress in single-vehicle perception, leveraging cameras [21], [22], LiDAR [23][24], radar [25], [26], and sensor fusion [27][28][29], [30], [31]. However, these systems are constrained by limited field of view, sensor blind spots, and occlusions, making them insufficient in complex, dynamic traffic scenarios.
@@
To address these limitations, Collaborative Perception (CP) leverages Vehicle-to-Everything (V2X) communication to enable real-time data sharing among vehicles and infrastructure [32], thereby enhancing the perception of occluded and long-range regions, improving safety, and expanding the Operational Design Domain [33], [34]. For example, at urban intersections, a vehicle may fail to detect occluded agents; V2X links allow perception data exchange with infrastructure or other vehicles to overcome such blind spots. V2X includes Vehicle-to-Vehicle (V2V), Vehicle-to-Infrastructure (V2I), Infrastructure-to-Infrastructure (I2I), and Vehicle-to-Network modes, enabling more comprehensive environmental awareness.
High-quality datasets are critical for algorithm development and benchmarking in CP [35]. Traditional datasets like KITTI [36], nuScenes [37], OpenMPD [38], and Waymo [2] focus on single-agent perception and lack support for inter-agent communication, making them unsuitable for CP research. Recently, numerous CP datasets, as illustrated in Fig. 1, have been released to support various scenarios, tasks, and sensor modalities, including DAIR-V2X [35], OPV2V [39], V2X-Sim [40], and V2X-Seq [41]. These datasets vary in scale, annotation quality, sensor layout, and latency modeling. A systematic review is necessary to clarify their characteristics, limitations, and application scope, and to inform future research directions in collaborative perception.
@@ I-A Roadmap
Figure 2: Chronological overview of collaborative perception datasets, including roadside, V2V, V2I, V2X, and I2I datasets.
Fig. 2 summarizes the chronological evolution of CP datasets across major collaboration paradigms. Development has progressed from static roadside perception to dynamic multi-agent collaboration, driven by increasing task complexity, richer sensors, and advanced communication technologies. Initial datasets, such as Ko-PER [42], CityFlow [43], and INTERACTION [44], focused on roadside perception without inter-agent communication. Around 2020, datasets like T&J [45], V2V-Sim [46], CoopInf [47], and CODD [48] introduced V2V and V2I collaboration, though often under constrained settings. Later efforts expanded scenario coverage: Adver-City [49] targets adverse weather, DeepAccident [50] supports motion and accident prediction, and H-V2X [51] covers highway scenes.
Since 2022, the focus has shifted to real-world deployment and sensor diversity. Datasets such as V2V4Real [52], V2X-Real [53], and DAIR-V2X-C [35] emphasize realism, while V2X-Seq [41], V2XReID [54], and DAIR-V2X-Traj [55] extend support to trajectory and identity tasks. V2X-Radar [56] and V2X-R [57] incorporate 4D radar, and V2XSet-Noise [58] simulates deployment noise for robustness benchmarking. Recent I2I datasets such as InScope [59] and RCooper [60] further diversify collaboration modes. Overall, CP datasets are evolving toward higher realism, richer modalities, and more specialized tasks.
@@ I-B Comparison to Related Reviews
This review distinguishes itself from prior reviews [61], [62], [63], [64], [65], [66], [67], [68], [69], [70], [71], which primarily focus on vehicle-centric datasets, specific techniques, or particular fusion strategies. While a few works partially overlap with ours in dataset coverage [72], [73], they overlook several recent CP datasets and benchmarks, and often lack systematic categorization or comparative depth. In contrast, our review provides a comprehensive and up-to-date compilation of CP datasets, with detailed analysis of their diversity, limitations, and suitability across various CP scenarios.
@@ I-C Main Contributions
The key contributions are summarized as follows:
Comprehensive Dataset Review: To the best of our knowledge, this is the first dedicated review that focuses exclusively on CP datasets. We systematically categorize and comprehensively analyze the existing datasets based on collaboration paradigms, data sources, sensor modalities, and application scenarios.
Task-Centric Classification: We organize datasets by core perception tasks in AD, including 3D object detection, semantic segmentation, and object tracking and prediction. This task-oriented taxonomy aids researchers in selecting appropriate datasets for specific applications. Recent developments and newly released datasets are also included to ensure timeliness.
Comparative Analysis and Future Trends: We conduct multi-dimensional comparisons, highlight representative datasets, and discuss emerging challenges such as standardization, LLM-based annotation, and cross-domain diversity, providing insights for future CP research.
@@ I-D Review Structure
The organization of this paper is illustrated in Fig. 3. Section II introduces CP datasets grouped by collaboration paradigms. Section III compares simulated and real-world datasets. Section IV focuses on application scenarios, while Section V analyzes sensor modalities. Section VI covers core perception tasks and evaluation metrics. Section VII provides dataset-level comparisons. Section VIII discusses key challenges and trends. Section IX concludes the paper. A summary of abbreviations is provided in Table I.
@@ II Collaboration-Based Perception Datasets
With the advancement of Intelligent Connected Vehicles (ICVs) and Intelligent Transportation Systems (ITS), CP paradigms have expanded beyond roadside perception to include V2V, V2I, V2X, and the emerging I2I frameworks. As illustrated in Fig. 4, these paradigms support diverse communication and fusion architectures. This section categorizes CP datasets by collaboration type: roadside (Section II-A), V2V (Section II-B), V2I (Section II-C), V2X (Section II-D), and I2I (Section II-E). Table II summarizes key datasets in each category. These resources are critical to driving progress in cooperative intelligence for next-generation ICVs and ITS.
@@
Griffin [95] 2025 arXiv C, L Sim 3DOD, MOT 275k - - Link
@@ II-A Roadside Datasets
Roadside perception plays a vital role in AD and ITS by offering a global, occlusion-resilient view of traffic participants [96][97][98]. Unlike vehicle-mounted systems, infrastructure-based perception units—typically installed on poles, gantries, or overpasses—enable the monitoring of complex intersections, dense traffic flows, and long-range environments with enhanced spatial stability and temporal continuity. By supplementing on-vehicle perception with infrastructure-level observations, roadside perception significantly improves perception robustness in scenarios involving occlusions, blind spots, or congested conditions.
Figure 5: Samples of roadside perception datasets. (a), (d) Ko-PER [42]; (b), (c), (e), (f) IPS300+ [75]; (g), (h) Rope3D [76]; (i) INTERACTION [44].
A variety of representative roadside datasets have been introduced to support infrastructure-based perception research. Early datasets such as Ko-PER [42] and CityFlow [43] primarily target intersections and Re-ID tasks. Later datasets expand coverage and functionality: INTERACTION [44] emphasizes interactive behaviors using drone-captured trajectories and semantic maps; Rope3D [76] provides top-down monocular images with 3D annotations; LUMPI [77] and IPS300+ [75] support multi-view, multi-modal fusion with rich annotations; and TUMTraf-I [78], A9-Dataset [74], RoScenes [79], and H-V2X [51] extend coverage to diverse weather, lighting, and highway scenarios.
Fig. 5 illustrates the diversity across Ko-PER [42], IPS300+ [75], Rope3D [76], and INTERACTION [44] in sensor modalities, viewpoints, and scene complexities. Ko-PER and Rope3D use monocular or stereo RGB with fixed viewpoints; Rope3D emphasizes long-range top-down detection, while Ko-PER captures lateral mid-range views. IPS300+ integrates LiDAR and stereo cameras for dense, occlusion-resilient 3D annotations. INTERACTION offers behavior-rich drone views for planning and imitation learning tasks. Together, they reflect trade-offs between perception complexity, spatial coverage, and task specialization.
These datasets have expanded infrastructure-centric perception research across detection, segmentation, and tracking tasks, while also enhancing V2I systems through broader scene coverage and perceptual redundancy. Nonetheless, challenges persist, including deployment costs, urban-centric biases, and difficulties in sensor synchronization and alignment across deployments.
@@ II-B V2V Datasets
V2V CP enhances perception coverage and accuracy by enabling vehicles to share information across viewpoints. It helps overcome occlusions and sensor limitations, supporting robust detection, semantic understanding, and motion prediction in multi-agent environments. High-quality datasets are thus vital for advancing V2V perception research.
V2V datasets have evolved significantly in alignment strategies and task diversity. Early datasets like OPV2V [39], LUCOOP [82], and MARS [83] assume ideal synchronization via GPS or timestamps, overlooking practical challenges such as asynchrony and drift. Later datasets incorporate asynchronous modeling (HEAL [84]), neural alignment (OPV2V-N [99]), and SLAM-based fusion (COMAP [80]). OPV2V+ [81] adds view selection and perception completion. While most support 2–4 vehicles, MARS emphasizes trajectory diversity. Tasks now cover detection, tracking, Bird’s Eye View (BEV) segmentation, and planning. V2V-QA [85] further introduces multi-modal reasoning using LLMs; IRV2V [100] and CODD [48] focus on registration under asynchrony.
Fig. 6 (a–c) illustrates this progression. OPV2V (a) provides synthetic CARLA data for controlled and reproducible benchmarking but lacks real-world complexity. V2V4Real (b) contributes large-scale real LiDAR data involving dual agents in diverse scenarios. V2V-QA (c) extends this by introducing multi-modal reasoning tasks powered by LLMs. Collectively, these datasets mark a shift from basic sensor fusion toward semantically rich and cognitively informed V2V collaboration.
Despite these advances, most V2V datasets still rely on offline stitching and neglect realistic modeling of communication delay, bandwidth constraints, and uncertainty. Future efforts should integrate real-time communication, dynamic alignment, and large-scale multi-agent urban scenes to support deployable cooperative strategies.
@@ II-C V2I Datasets
V2I perception datasets enhance AD systems by facilitating collaboration between vehicles and roadside infrastructure. These datasets integrate multimodal information from both vehicle-mounted and infrastructure-based sensors, enabling a more accurate and holistic understanding of the driving environment.
Representative V2I datasets exhibit distinct characteristics and collaboration designs. CoopInf [47] utilizes eight infrastructure sensors to support multi-view 3D object detection. DAIR-V2X-Seq [41] is the first large-scale sequential V2X dataset incorporating time-aligned data from both vehicles and infrastructure. HoloVIC [86] creates holographic intersection scenes with diverse layouts, aiming to mitigate occlusion and blind spots through multimodal fusion. TUMTraf-V2X [88] focuses on perception in challenging lighting conditions and supports 3D detection and tracking tasks. OTVIC [87] addresses high-speed and noisy communication settings using multimodal, multi-view sensor data.
Figure 6: Samples from representative V2V and V2I datasets. (a) OPV2V [39]; (b) V2V4Real [52]; (c) V2V-QA [85]; (d)(e)(g)(h) DAIR-V2X [35]; (f)(i) OTVIC [87].
As illustrated in Fig. 6 (d)–(i), DAIR-V2X ((d)(e)(g)(h)) provides synchronized vehicle-side and infrastructure-side data across various lighting and weather conditions, making it suitable for robust V2I perception under real-world deployment constraints. In contrast, OTVIC ((f)(i)) focuses on sparse highway scenarios with long-range detection tasks, leveraging BEV projections and trajectory-level annotations. The visual samples highlight the contrast in scene layout, detection distance, and data density, reflecting complementary application targets.
While these datasets significantly contribute to improving V2I CP, challenges remain. Sensor synchronization is technically demanding, and many datasets offer limited scene diversity or weather coverage [101]. Future research should prioritize more heterogeneous scene types and realistic multi-view synchronization to enhance generalization across V2I deployments.